OpenStudy (anonymous):

HELP!!!!!!!!!! Find the largest number which is a factor of the numbers 130,169 and 195....

OpenStudy (anonymous):

use the number rules

OpenStudy (anonymous):

which number rule

OpenStudy (anonymous):

i dnt kno any number rule

OpenStudy (anonymous):

things like if it ends in an even number its divisible by 2, if you add all the numbers up and its a multiple of 3 its divisable by 3, if it ends in a 0 or 5 its divisable by 5, if it ends in 0 its divisable by 10

OpenStudy (anonymous):

13 is the largest common factor 130/13 = 10 169/13 = 13 195/13 = 15

OpenStudy (anonymous):

thnx @surdawi